a penny is tossed and a number cube is rolled. find the probabilities.
1. P(tails, 4)

2. P(heads, odd number)

3.P(heads, not three)

4. P(tails, 7)

Answer by venugopalramana(3286)

a penny is tossed and a number cube is rolled. find the probabilities.
1. P(tails, 4)
2. P(heads, odd number)
3.P(heads, not three)
4. P(tails, 7)
IN TOSSING OF A PENNY,THERE ARE ONLY 2 POSSIBILE OUTCOMES HEADS AND TAILS...SO P(TAILS)=P(HEADS)=1/2
IN THROWING OF CUBE THERE ARE 6 POSSIBLE OUTCOMES.SO PROBABILITY ANY ONE NUMBER TURNING UP IS 1/6
THESE 2 ACTS OF THROWING A PENNY AND THROWING A CUBE ARE INDEPENDENT EVENTS...SO THEIR COMPOUND PROBABILTY IS GIVEN BY THE PRODUCT OF THE 2 PROBABILITIES.....SO....
1. P(tails, 4).....1/2*1/6=1/12
2. P(heads, odd number) ...THERE ARE 3 ODD NUMBERS..1,3 AND 5..SO PROBABILITY OF ODD NUMBER TURNING UP IS 3/6=1/2
SO THE COMPOUND PROBAILTY IS 1/2*1/2=1/4
3.P(heads, not three) ......NOT 3 MEANS 5 POSSIBILITIES 1,2,4,5 AND 6 SO PROBABILITY IS 5/6
SO COMPOUND PROBABILIY IS 1/2*5/6=5/12
4. P(tails, 7)......1/2*1/6=1/12